# Quantifying and Mitigating Unimodal Biases in Multimodal Large Language Models: A Causal Perspective

### v3

@@ 3.3 Multiple-Choice Question Construction
In this subsection, we detail the process of constructing multiple-choice questions with four candidate options.
@@
After obtaining the subgraph that meets the criteria, we generate questions using the entity-relation path in it. Specifically, to obtain fluent and coherent questions, we feed the path into an LLM to produce the target question text. Our inputs include the names of start entity and all relations within the path. We adopt the in-context learning (ICL) technique Brown et al. (2020) and provide several examples to LLMs. The prompt template is in Appendix A.1. Comparing various LLMs with adjusted instructions, we find ChatGPT generates the highest-quality multi-hop questions, hence choosing its results for subsequent human evaluation. Finally, to prevent information leakage, entity names in the question are replaced with “this <OBJECT_NAME>”.
@@
As discussed in Section 2.2, language bias refers to the model overly attending to the information from questions through the undesired path ${Q}\rightarrow S\rightarrow A$ or ${Q}\rightarrow T\rightarrow A$. To simulate such a situation, we query an MLLM with the generated question in a text-only setup. To ensure that the final options are identical for all MLLMs, we uniformly use the answers obtained from querying GPT-4V. The output for the mentioned example in Figure 3 is “Tesla”. The prompt template we use can be seen in Appendix A.2.
@@
To explore vision bias along the path $I\rightarrow A$, we include the visually associated entity name (e.g., “Chrysler PT Cruiser”) as an option. This allows us to observe whether the model directly selects it upon encountering an option that aligns with the visual information.
@@
Moreover, we introduce a semantic misleading option, such as “The Chrysler Corporation”, to challenge the multi-hop reasoning of MLLMs in KG. This option refers to the entity that is pointed by the relation commonly owned by both the associated entity and its sampled neighbor. For example in Figure 3, upon encountering a question about “brand” and “Chrysler PT Cruiser”, MLLMs might simply output a direct answer (e.g., “The Chrysler Corporation”), ignoring other constraints in the question (e.g., “followed by”), hence struggling to choose the correct answer (e.g., “Fiat Automobiles S.p.A.”).
@@
Corresponding to the causal path via ${I}\rightarrow G$ and ${S}\rightarrow G$, this option is the end entity of the entity-relation path (e.g., “Fiat Automobiles S.p.A.”). Finally, we check and ensure that each option is distinct from the remaining three to eliminate any potential overlap in samples.
3.4 Causal Rationale Generation
Furthermore, the entity-relation paths can assist in generating the reasoning rationale, termed as causal rationale, tailored to the question at hand. In this context, we have employed a heuristic rule-based approach, starting from the associated entity and generating a causal rationale step by step until reaching the ground truth answer. Templates of generation can be found in the Appendix A.3. These generated causal rationales can be utilized to verify whether the reasoning process of MLLMs is correct, thereby providing interpretability. They can also be used to fine-tune specialized MLLMs to enhance their multi-hop reasoning capabilities, with experimental results available in Section 5.3. The causal rationales could further be polished and refined through LLMs such as ChatGPT, and we leave this for future work.
3.5 Dataset Statistics
@@
Using the aforementioned construction approach, we automatically generate training data from the train set of INFOSEEK, and development/test data from the validation set of INFOSEEK. Table 4 in Appendix shows the statistics of our curated dataset MORE. For the purpose of generating questions with depths of 2-hop and 3-hop, we set $n=1,2$, respectively. We do not include longer-hop queries here due to concerns that their complexity will render the generated questions ambiguous or difficult to comprehend.
@@ 3.6 Quality Analysis
To ensure the quality of the comprising datasets, we analyze the lexical diversity and the fluency of the generated questions. The considered baselines and metrics are detailed in Appendix C.1. From Figure  4, MORE shows superiority in terms of lexical diversity and fluency, even in comparison to datasets created through human annotator-based question generation.
@@
We also conduct a human evaluation to validate and assess the quality of the generated questions. Results in Appendix C.2 show that 90.6% of the generated questions are classified as valid by the annotators, further demonstrating the quality of our datasets.
4 Evaluating MLLMs on MORE
Model LLM # Params MORE (Two-Hop, acc (%)) MORE (Three-Hop), acc (%) MORE (Overall, acc (%))
Open-ended Multi-choice Open-ended Multi-choice Open-ended Multi-choice
Random / / / 25.0 / 25.0 / 25.0
BLIP-2 OPT 6.7B 4.0 16.4 1.4 15.4 2.7 15.9
InstructBlip Vicuna 13B 3.0 17.0 1.6 16.2 2.3 16.6
mPLUG-Owl Llama 7B 4.0 12.4 8.2 11.4 6.1 11.9
LLaVA Llama 13B 8.0 20.8 6.8 13.6 7.4 17.5
GPT-4V - - 15.8 25.6 15.3 23.2 15.6 24.4
Gemini Pro - - 14.2 33.5 10.1 24.4 12.2 28.9
Table 2: MLLMs’ results on the test set of MORE. We report the VQA accuracy (%) under the open-ended generation and the multi-choice settings on two-hop, three-hop, and all data, respectively.“-” denotes not released information.
@@ 4.1 Experimental Setup
We use all the test data of MORE for evaluation. We adopt two different settings: 1) Open-ended. Ask the MLLM to generate answers based on the input image and question, or 2) Multi-choice. Provide MLLM with four options and let it choose the correct answer from them. The latter setting has a random baseline (accuracy of 25%).
@@
We evaluate various leading MLLMs on our MORE dataset in a zero-shot fashion, including two limited-access MLLMs: GPT-4V OpenAI (2023) and Gemini Pro Team et al. (2023), and four open-source MLLMs: BLIP-2 (6.7B)  Li et al. (2023), InstructBLIP (13B)  Dai et al. (2023), mPLUG-Owl (7B) Ye et al. (2023), and LLaVA (v1.5, 13B) Liu et al. (2023a), detailed in Appendix D.1. As for evaluation metrics, we adopt the VQA accuracy Antol et al. (2015) for all models for a fair comparison. We also quantify the causal effects of images and questions on models’ predictions in Section 4.3.
@@ 4.2 Evaluation Results
We present the results of MLLMs separately on two-hop, three-hop, and all data of MORE in Table 2. Examples of model prediction can be found in Appendix H. We observe that: 1) The performance of all baselines on MORE is poor (e.g., under the “Multi-choice” setting, only Gemini Pro exceeded the random baseline at 28.9% accuracy), indicating MLLMs’ vulnerability to language and vision biases. 2) There is still a gap between open-source and limited-access models on MORE, especially under the “Open-ended” setting. 3) Most models perform better on two-hop data than on three-hop data (Gemini Pro is particularly outstanding on two-hop data, achieving 33.5% accuracy), suggesting that MLLMs’ reasoning capabilities are challenged when the problems become more complex. 4) GPT-4V performs best under the “Open-ended” setting, but falls short under the “Multi-choice” setting versus Gemini Pro, possibly due to that we use homologous ChatGPT-generated distractors when constructing the language bias option in Section 3.3, which poses a greater challenge to GPT-4V’s judgment. This point is also verified in the following analyses in Section 4.3.
@@ 4.3 Causal Analysis of VQA Biases
In this subsection, we analyze the performance of MLLMs through a causal lens.
@@
In Figrue 5, we show the distribution of options under the “Multi-choice” setting across various MLLMs. We observe that: 1) BLIP2 and GPT-4V often incorrectly choose options indicating language bias, aligning with our prior analysis on GPT-4V. 2) The proportion of either language or vision bias exceeded 40% across all the models, demonstrating the significant impact of unimodal biases on their predictions. 3) To some extent, models’ selection of semantically misleading options indicates some ability to combine visual and textual information, though not fully grasping the problem. This highlights the challenge our MORE dataset poses to current MLLMs. Please note that discrepancies may exist between the proportions of ground truth options presented here and the accuracy values reported in Table 2, as some models’ outputs (e.g., mPLUG-Owl) may not align with the provided options, thus affecting the count of valid answers.
@@
To further analyze the impact of vision bias and language bias on the model’s prediction, we assess the causal effects based on the definitions provided in Section 2.3. Specifically, we randomly select 100 samples for intervention and then measure the average of the effects over all instances to calculate $\operatorname{TCE}$ (corresponds to model’s sensitivity) and $\operatorname{DCE}$ (corresponds to models’ robustness). We implement two straightforward and natural intervention ways separately for questions and images. For questions: (i) we change both the core semantics and the textual form, or (ii) we only change the textual form while controlling to keep the core semantics unchanged. For images: (i) we change both the entities of interest and the visual context, or (ii) we only change the visual context while keeping the core entities unchanged. As discussed below Equation (9), we do not conduct sole interventions on the core entity while concurrently controlling for the irrelevant visual context, the ground truth answer, and the question. This is because it is extremely difficult to find such qualified samples to calculate $\operatorname{DCE}({E}\rightarrow A)$ in practical scenarios. Overall, a higher $\operatorname{TCE}$ is desirable, indicating better sensitivity, while a lower $\operatorname{DCE}$ indicates better robustness. Implementation details of the intervention can be found in Appendix E.
From Figure 6: 1) Current MLLMs exhibit high sensitivity (high $\operatorname{TCE}$), a possible reason is that instruction tuning makes models sensitive to variations in input Stolfo et al. (2022). 2) However, the robustness is relatively low (high $\operatorname{DCE}$), showing predictions change with variations in input even when the ground truth is fixed, suggesting the reliance on spurious paths over genuine causal features. 3) The relationship between MLLMs’ utilization of biases and their performance is intricate, with GPT-4V performing well even amid strong language and vision biases.
5 Mitigating VQA Biases for MLLMs
Model / Acc (%) OK-VQA INFOSEEK MORE
Open-ended. Multi-choice.
GPT-4V 70.4 38.0 15.6 24.4
Gemini Pro 64.9 37.8 12.2 28.9
LLaVA 72.8 16.0 7.4 17.5
GPT-4V (w. the DeVA framework) 74.8 50.6 24.2 30.3
Gemini Pro (w. the DeVA framework) 68.3 50.3 19.5 32.1
LLaVA (trained on INFOSEEK) 72.1 19.7 3.1 13.9
LLaVA (trained on MORE) 73.6 20.2 16.4 28.9
Table 3: Results comparison after employing our strategies on OK-VQA, INFOSEEK, and MORE.
In this section, we explore two strategies to mitigate the unimodal biases of MLLMs.
5.1 Experimental Setup
Dataset
We conduct experiments on three knowledge-based VQA datasets OK-VQA Marino et al. (2019), INFOSEEK Chen et al. (2023), and our curated MORE, where relying solely on image content proved insufficient for answering questions. We randomly select 1,000 instances from the validation set of INFOSEEK and OK-VQA, respectively, and report results on the test data of MORE, to ensure a consistent number of samples for evaluation. Since INFOSEEK and OK-VQA do not provide multiple options, we only report the results of them under the “Open-ended” settings.
Baselines
We focus on three models that have shown superior performance in Table 2: LLaVA, GPT-4V, and Gemini Pro. For limited-access models GPT-4V and Gemini Pro, we propose a prompting-based Decompose-Verify-Answer (DeVA) framework. For the open-source model LLaVA, we conduct fine-tuning on various datasets and compare the results. We do not apply the DeVA framework to LLaVA based on findings from our pilot study that LLaVA’s performance in comprehending and responding to our instructions is not on par with that of GPT-4V and Gemini Pro.
5.2 A Decompose-Verify-Answer Framework
For a given instance, DeVA begins with a question decomposer to break down the complex question into several easier ones step-by-step, so as to explicitly avoid the model simply taking a spurious path to give the answer. For each subquestion, after providing an answer, DeVA employs a verifier to confirm the accuracy of the answer. The verifier strategizes the utilization of external tools and investigates their outputs, thereby acquiring the necessary context or information needed to provide a precise answer, such as image retrieval and text retrieval. This iterative process of answering and verifying will continue until all subquestions have been resolved. By incorporating external knowledge, the final verified output comprehensively integrated the information of the whole reasoning process to give a correct answer. A detailed illustration of DeVA and the prompt template is in Appendix F.
Results
From Table 3, we can see that: 1) the DeVA framework significantly enhances the performance of GPT-4V and Gemini Pro across all three datasets. 2) Notably, the improvement on the INFOSEEK dataset is more pronounced than on MORE. This is because INFOSEEK is a dataset focused on information-seeking, and once we provide the models with relevant external knowledge, they can quickly extract the necessary information and output a correct answer. However, the limited results on MORE suggest that even when models are allowed access to additional information, they may still be impeded by language and vision biases. Therefore, how to disentangle and address this issue remains a challenging problem, which we leave for future work.
5.3 Fine-tuning LLaVA
We fine-tune LLaVA using the LoRA Hu et al. (2022) technique on the training sets of INFOSEEK and MORE, respectively. For MORE, we incorporate the generated causal rationale into the instruction of some samples, so as to enhance the models’ reasoning ability. Other implementation details are in Appendix G. From Table 3, we can see: 1) After fine-tuning, LLaVA shows a significant improvement in performance on OK-VQA, INFOSEEK, and MORE, particularly LLaVA (trained on MORE) outperforms the vanilla GPT-4V on OK-VQA and MORE. 2) While LLaVA (trained on INFOSEEK) does improve its performance on INFOSEEK, it significantly reduces its effectiveness on MORE. In contrast, LLaVA (trained on MORE) demonstrates better generalizability, confirming the effectiveness of incorporating rationales into instructions.
@@ 6 Related Work
Multimodal Large Language Models
Recent advancements in LLMs have facilitated the development of powerful MLLMs. Pioneer work has demonstrated their impressive capability on multimodal tasks OpenAI (2023); Team et al. (2023); Liu et al. (2023a). Despite the substantial success, the current evaluation of MLLMs mainly focuses on basic visual tasks Liu et al. (2023b); Fu et al. (2023); Lu et al. (2024), and their reasoning ability has not been investigated in detail.
@@
Several previous works have studied knowledge-based VQA problems Wang et al. (2017); Marino et al. (2019); Chen et al. (2023). While these datasets do require knowledge to answer questions, they have the following limitations: 1) They often focus solely on image-related information without requiring the model to perform multi-hop reasoning. 2) The answer types are typically open-ended and do not provide reasoning rationales, making it difficult to evaluate the model’s output. 3) They cannot assess the degree of reliance on language or visual biases by models. As initially introduced in Section 1, Table 1 provides a comparative summary of various datasets alongside our MORE dataset, underscoring its enhanced comprehensiveness.
@@
Studies have found that some VQA models tend to rely on statistical priors present in the training data, rather than truly understanding the content of images Agrawal et al. (2018). This issue manifests in two primary forms: language bias and vision bias. The former emerges from strong correlations between specific questions and their answers Abbasnejad et al. (2020); Zhu et al. (2020), and the latter stems from the frequent co-occurrence of certain textual and visual elements in the training dataset Si et al. (2022); Gupta et al. (2022). Recent strategies to mitigate such biases mainly focus on employing data augmentation techniques Niu et al. (2021). Although the concerns are similar, all the aforementioned methods focus on the statistical bias under the pre-train and fine-tune paradigm, while we focus on the reasoning ability of MLLMs.

### v4

@@ 3.2 Multiple-Choice Question Construction
We detail the multiple-choice questions construction with four options, guided by the unimodal bias definition in Section 2.2.
@@
We generate questions by analyzing entity-relation paths within a specified subgraph, converting these paths into fluent, coherent queries with an LLM. We utilize the in-context learning (ICL) technique Brown et al. (2020) and a standardized prompt (shown in Appendix A.1) to ensure quality. Among various LLMs tested, ChatGPT was selected for its superior multi-hop question generation. To avoid information leakage, entity names in the question are anonymized as “this <ENTITY_NAME>”. For instance, a final generated question asks “What brand follows this vehicle in its product line?” in Figure 3.
@@
Language bias occurs when models overly attend to question-related information via pathways like ${Q}\rightarrow T\rightarrow A$. To assess this, we simulate conditions where only the question is provided and evaluate MLLMs’ responses. We use the answers of GPT-4 (i.e., the text-only version of GPT-4V), to ensure uniform final options across MLLMs. For example, in Figure 3, GPT-4 reponds “Tesla” to the question. Additional results using different models for generating language bias options are detailed in Section 5.3 and the prompt template is available in Appendix A.2.
@@
Vision bias occurs when visual information dominates (e.g., via $I\rightarrow{E}\rightarrow A$). We use the visually associated entity name (e.g., “Chrysler PT Cruiser”) as an option, to see if the model directly selects it upon encountering an option that aligns with the visual information.
@@
We introduce a semantic misleading option, such as “The Chrysler Corporation”, to challenge MLLMs’ multi-hop reasoning. This option refers to the entity that is pointed by the relation commonly owned by both the associated entity and its sampled neighbor. For example in Figure 3, upon encountering a question containing “brand” and “Chrysler PT Cruiser”, MLLMs might simply output a direct answer (e.g., “The Chrysler Corporation”), ignoring other constraints in the question (e.g., “followed by”), hence struggling to choose the correct answer (e.g., “Fiat Automobiles S.p.A.”).
@@
Corresponding to the causal path via ${E}\rightarrow G$ and ${S}\rightarrow G$, this option represents the final entity in the entity-relation path (e.g., “Fiat Automobiles S.p.A.”). Finally, we check and ensure that each option is unique to prevent overlap samples.
Causal Rationale Generation
Using entity-relation paths, we generate a causal rationale that aids in answering questions through a heuristic rule-based approach. As shown in Figure 3, this process begins with the associated entity and proceeds step-by-step to the ground truth. These rationales help confirm the accuracy of MLLMs’ reasoning and improve their interpretability. They also can contribute to fine-tuning MLLMs for better multi-hop reasoning.
3.3 Dataset Statistics and Quality Analysis
@@
We automatically generate training data from INFOSEEK’s train set, and development/test data from its validation set, as shown in Table 3 in Appendix. Focusing on questions with 2-hop and 3-hop depths, we set $n=1,2$ respectively. We avoid longer-hop questions to prevent potential ambiguity and complexity.
@@
We analyze the lexical diversity and fluency of the generated questions, with baselines and metrics detailed in Appendix C.1. From Figure 4, MORE shows superiority in lexical diversity and fluency, even compared to human-generated datasets.
@@
Our human evaluation confirms the high quality of generated questions and rationales, with 91% of questions and 98% of rationales deemed valid by annotators (details are in Appendix C.2).
3.4 Causal Effect Calculation on MORE
Based on the analysis in Section 2.3, we discuss how to quantify the causal effects of unimodal biases on model predictions using MORE.
Causal Effect of Questions
When intervening on questions, we keep the images constant: (1) For $\operatorname{TCE}({Q}\text{ on }A)$, we alter the question to another from the dataset pertaining to the same entity, changing both its text and core semantics, which shifts the ground truth. (2) For $\operatorname{DCE}({T}\rightarrow A)$, we let ChatGPT rephrase the question, altering its textual form but preserving its semantic meaning, resulting in no ground truth change post-intervention.
Causal Effect of Images
When intervening on images, we keep the questions constant: (1) For $\operatorname{TCE}({I}\text{ on }A)$, we replace the image with another from the dataset corresponding to the same question but featuring different entities, altering both core entity and visual context to change the ground truth post-intervention. (2) For $\operatorname{DCE}({C}\rightarrow A)$, we replace the image with another that depicts the same entity, to keep the ground truth consistent after the intervention. Besides, it is practically challenging to calculate $\operatorname{DCE}({E}\rightarrow A)$. Because intervening on core entity $E$ while maintaining the constancy of the question $Q$, visual context $C$, and ground truth answer $G$ is hard to achieve, hence not considered in further analysis.
Given the impracticality of testing all perturbations of $T$ and $C$, we randomly select 100 samples for each type of intervention and compute the average effects to determine $\operatorname{TCE}$ and $\operatorname{DCE}$ in Section 5.3. Overall, a higher $\operatorname{TCE}$ indicates better sensitivity, while a lower $\operatorname{DCE}$ indicates better robustness.
4 CAVE for Bias Mitigation
To mitigate unimodal biases in MLLMs and improve their reasoning capabilities, we propose CAVE, a causality-enhanced agent framework in this section.
For a given instance, CAVE starts by using a question decomposer to break down complex questions into simpler, step-by-step subquestions, explicitly avoiding spurious paths that may lead to incorrect answers. For each subquestion, CAVE uses a causality-enhanced reasoner to evaluate the correctness of the decomposed subquestion. Specifically, based on the aforementioned causal analysis in Section 2, it assesses whether the decomposed subquestion changes under two conditions: 1) rephrasing the question text, or 2) replacing the image with another captured by the same entity from a different angle or at a different time. If it identifies that the decomposed subquestion needs to be altered, this indicates that the previous understanding is influenced by irrelevant factors or biases, failing to capture the true semantics. In such cases, it will enforce a new round of question decomposition until it determines that the current subquestions are appropriate and accurate. Next, CAVE uses a verifier to strategically employ external tools and evaluate their outputs, acquiring the necessary context or information, such as image and text retrieval, to provide a precise answer. This iterative process of answering and verifying continues until all subquestions are resolved. By incorporating external knowledge, the final verified output integrates information from the entire reasoning process to provide a correct answer. A detailed illustration of CAVE and the prompt template is provided in Appendix E.
5 Experiments
Model LLM # Params Two-Hop, acc (%) Three-Hop, acc (%) Overall, acc (%)
Random / / 25.0 25.0 25.0
InstructBlip Vicuna 13B 17.0 16.2 16.6
mPLUG-Owl Llama 7B 12.4 11.4 11.9
LLaVA Llama 13B 20.8 13.6 17.5
Qwen-VL Qwen 7B 17.4 15.6 16.5
GPT-4v - - 17.3 16.0 16.5
GPT-4o - - 18.0 17.0 17.5
GPT-4o + CAVE - - 33.7 27.3 30.5
Gemini Pro (V) - - 25.4 20.7 22.3
Gemini Pro (V) + CAVE - - 35.6 28.8 33.2
Table 2: MLLMs’ results on the test set of MORE. We report the VQA accuracy (%) under the multi-choice settings on two-hop, three-hop, and all data, respectively.“-” denotes not released information.
@@ 5.1 Experimental Setup
We evaluate all test data from the MORE dataset using the Multi-choice settings, where MLLMs select answers from four provided options, with a random baseline accuracy of 25%.
@@
We evaluate various leading MLLMs on our MORE dataset in a zero-shot fashion, including three limited-access MLLMs: GPT-4v, GPT-4o OpenAI (2023), and Gemini Pro Vision Team et al. (2023), and four open-source MLLMs: InstructBLIP (13B)  Dai et al. (2023), mPLUG-Owl (7B) Ye et al. (2023), LLaVA (v1.5, 13B) Liu et al. (2023a), and Qwen-VL (7B) Bai et al. (2023b) the details are in Appendix D. For consistent evaluation, we use standard accuracy metrics for all the models. We further quantify the causal effects of images and questions on model predictions following Section 5.3.
@@ 5.2 Evaluation Results
The results of MLLMs on MORE are shown in Table 2 and further exemplified in Appendix F. We observe that:
1) All baselines perform poorly on MORE (e.g., even the best-performing model, Gemini Pro Vision, achieves only 22.3% accuracy, which does not surpass the random baseline), indicating MLLMs’ vulnerability to biases.
2) Most models perform better on two-hop data than on three-hop data, suggesting that MLLMs’ reasoning capabilities are challenged when the problems become more complex.
3) GPT-4v falls short versus Gemini Pro Vision, possibly because we use homologous GPT-generated distractors when constructing the language bias options in Section 3.2, which particularly challenges GPT-4v’s judgment. This point is further analyzed in Section 5.3.
4) Our proposed CAVE significantly enhances GPT-4o and Gemini Pro Vision on MORE, validating the effectiveness of our method. However, the relatively low absolute values indicate ongoing challenges related to biases, suggesting the need for further research efforts.
@@ 5.3 Causal Analysis of VQA Biases
In this subsection, we select several representative MLLMs and analyze their performance through a causal lens.
@@
In Figure 5, we show the option distribution of selected MLLMs. Here we mainly utilize language bias options generated by GPT-4v and Gemini Pro Vision for comparison because they explicitly provide corresponding text-only versions. The observations are as follows:
1) Severe unimodal biases. More than 40% of the options show either language or vision bias in all models, underscoring the prevalence of unimodal biases.
2) Limited understanding ability. Models’ selection of semantically misleading options indicates some ability to combine visual and textual information, though not fully grasping the problem. This highlights the challenge our MORE dataset poses to current MLLMs.
3) Obvious selection tendency. GPT-4v often incorrectly chooses language bias options generated from GPT-4 (i.e., the text-only version of GPT-4v). Switching to Gemini Pro (i.e., the text-only version of Gemini Pro Vision) shifts this trend, with GPT-4v’s language bias selections decreasing and Gemini Pro’s increasing. These observations align with our prior analysis of GPT-4v.
Please note that discrepancies may exist between the proportions of ground truth options presented here and the accuracy values reported in Table 2, as some models’ outputs may not align with the provided option format (e.g., mPLUG-Owl), thus affecting the count of valid answers.
@@
To further analyze the impact of unimodal biases on the model predictions, we assess the causal effects based on the discussion in Section 3.4. As shown in Figure 6, we select several representative MLLMs:
1) Current MLLMs exhibit high sensitivity (high $\operatorname{TCE}$), a possible reason is that instruction tuning makes models sensitive to variations in input Stolfo et al. (2023).
2) However, the model’s robustness is relatively low, as indicated by a high $\operatorname{DCE}$, meaning its predictions fluctuate even when the ground truth remains constant. This phenomenon becomes particularly pronounced when irrelevant text surface forms are introduced. This suggests that, although MLLMs can adapt well to answers, they exhibit a lack of robustness when exposed to interference, relying on spurious correlations rather than genuine causal features.
@@ 6 Related Work
Multimodal Large Language Models (MLLMs)
Recent advances in large language models (LLMs) have led to the emergence of MLLMs, which demonstrate exceptional performance in multimodal tasks OpenAI (2023); Team et al. (2023); Liu et al. (2023a); Hu et al. (2023); Yu et al. (2024). However, in contrast to the extensive evaluation of reasoning capabilities in LLMs Wei et al. (2022); Chen et al. (2024a), currently, the evaluation of MLLMs primarily emphasizes basic visual tasks Liu et al. (2023b); Fu et al. (2023); Lu et al. (2024); Chen et al. (2024b), with limited investigation into their reasoning capabilities.
@@
Existing VQA datasets Wang et al. (2017); Marino et al. (2019); Chen et al. (2023) are limited by their focus on image-related information, lack of multi-hop reasoning, open-ended answers, and reasoning rationales. They also fail to measure the effect of language and vision biases. Our MORE dataset, outlined in Table 1, addresses these shortcomings by providing a more comprehensive assessment.
@@
Research reveals that some VQA models primarily depend on statistical priors from training data instead of genuinely comprehending image content Agrawal et al. (2018). These models exhibit language and vision biases; the former arises from strong correlations between specific questions and answers Abbasnejad et al. (2020); Zhu et al. (2020), and the latter from frequent co-occurrences of textual and visual elements in the dataset Si et al. (2022); Gupta et al. (2022). Recent efforts to address these biases mostly involve data augmentation Niu et al. (2021). Besides, several studies Rohrbach et al. (2018); Parcalabescu et al. (2022); Parcalabescu and Frank (2023) use examples derived from language biases to assess object hallucination in models. However, these constructed examples are typically counterfactual, with the answer types limited to yes/no, essentially judging whether the constructed statements are correct or incorrect.